OpenStudy (lgbasallote):

a steel ball is placed at time t=0 in a medium which is maintained at a temperature of 30 C. heat flows so rapidly within the ball that the temperature is essentially the same at all points of theball. at the end of 3 minutes, the temperature of the ball reduced tp 70 C. after 15 minutes the temperature of the ball becomes 42 C. find the initial temperature of the ball

OpenStudy (lgbasallote):

Given: T_m = 30 C @ t = 0 T = ? @ t = 3 T = 70 @ t = 15 T = 42

OpenStudy (lgbasallote):

\[\ln (\frac{42 - 30}{70-30} )= k(15-3)\] \[\ln (0.3) = 12k\] \[k= -0.100\]

OpenStudy (lgbasallote):

\[\ln (\frac{T - 30}{42 - 30}) = -0.100 (3)\] \[\ln (\frac{T - 30}{12}) = -0.301\] \[\frac{T-30}{12} = e^{-3.01}\] \[T = 30 + 12e^{-3.01}\] \[T = 38.89 C\] what did i do wrong

OpenStudy (anonymous):

−0.100(0-15) not −0.100(3)

OpenStudy (anonymous):

first line of the last reply

OpenStudy (anonymous):

\[\ln(\frac{T-30}{42-30})=-0.1(0-15)\]

OpenStudy (lgbasallote):

oh lol of course thanks

OpenStudy (anonymous):

yw

OpenStudy (lgbasallote):

so \[\ln (\frac{T-30}{12}) = 1.5\] \[T = 30 + 12e^{1.5}\] \[T = 84^o C \checkmark\]